# FaithAct: Faithfulness Planning and Acting in MLLMs

### v1

@@ Abstract
Unfaithfulness remains a persistent challenge for large language models (LLMs), which often produce plausible yet ungrounded reasoning chains that diverge from from either perceptual evidence or their final conclusions. We distinguish between behavioral faithfulness (alignment between reasoning and output) and perceptual faithfulness (alignment between reasoning and input), and introduce FaithEval for quantifying step-level and chain-level faithfulness by evaluating whether each claimed object is visually supported by the image. Building on these insights, we propose FaithAct, a faithfulness-first planning and acting framework that enforces evidential grounding at every reasoning step. Experiments across multiple reasoning benchmarks demonstrate that FaithAct improves perceptual faithfulness by up to 26% without degrading task accuracy compared to prompt-based and tool-augmented baselines. Our analysis shows that treating faithfulness as a guiding principle not only mitigates hallucination but also leads to more stable reasoning trajectories. This work thereby establishes a unified framework for both evaluating and enforcing faithfulness in multimodal reasoning.
FaithAct: Faithfulness Planning and Acting in MLLMs
Junxian Li1thanks: Equal contribution.   Xinyue Xu211footnotemark: 1   Sai Ma311footnotemark: 1   Sichao Li4thanks: Correspondence: scli@cityu.edu.mo 1Shanghai Jiao Tong University 2The Hong Kong University of Science and Technology 3The Australian National University 4City University of Macau
@@ 1 Introduction
Despite rapid progress in multimodal large language models (MLLMs) 2023llavarlhf; chen2024expanding; an2025llavaonevision15fullyopenframework; Bai2025Qwen25VLTR, their reasoning traces remain unfaithful: models frequently produce persuasive explanations that conflict with perceptual evidence, or utilize post-hoc stories pretending their reasoning progress arcuschin2025chain; barez2025chain. This gap poses a central challenge for trustworthy reasoning. Existing efforts typically focus on improving task accuracy or enriching chain-of-thought (CoT) generation zhang2023multimodal, yet they rarely ensure faithfulness as a principle.
@@
We are motivated by the following principle and observations goyal2017making.
@@
This principle echoes long-standing findings in Visual Question Answering (VQA): systems should avoid answering beyond available evidence and resist over-reliance on language priors agrawal2018don.
Figure 1: Examples of perceptual and behavioral unfaithfulness. Left: Behavioral unfaithful and perceptual unfaithful. The model presents a step-by-step reasoning trace describing alternating arrow directions and increasing dot counts, yet such reasoning does not reflect its actual decision process. The final choice (A) is likely made through pattern association, with the explanation generated post hoc to rationalize it. Right: Behavioral faithful but perceptual unfaithful. The model describes the bicycle as yellow, influenced by the nearby yellow bus, even though the bicycle is gray. This illustrates a visually plausible but perception-unfaithful reasoning step, where linguistic association overrides perceptual grounding.
As illustrated in Figure 1, MLLMs can produce seemingly correct answers and fluent explanations while exhibiting unfaithful reasoning. We highlight two failure modes: (i) a post-hoc narrative that does not reflect its decision process; and (ii) claims about objects (e.g., cars, arrows, and dots) that do not match the input.
We therefore disentangle behavioral faithfulness (reasoning steps align with the model’s output) and perceptual faithfulness (reasoning steps align with the model’s input) arcuschin2025chain; matton2025walk; ming2024faitheval; Li2023EvaluatingOH. We argue that faithfulness should be a design principle, not merely an evaluation objective. Reasoning frameworks should explicitly verify the evidential grounding of each step before it is admitted into the chain, ensuring that inference remains both perceptually justified and behaviorally aligned. This motivates FaithAct, a faithfulness-first planning framework that integrates evidence verification into every stage of reasoning, turning free-form chains of thought into chains of evidence.
@@
Our main contributions are:
@@
  • We introduce FaithEval, a general perceptual faithfulness evaluation pipeline that extracts claimed objects from each reasoning step and verifies their existence at both local and global levels through preference polling and visual grounding. For every step, FaithEval computes a step-level faithfulness score ($F_{\text{step},t}$) and aggregates them into a chain-level score ($F_{\text{chain}}$), providing a quantitative measure throughout the reasoning process.
  • We propose FaithAct, a faithfulness-first planning and acting framework that enforces evidential grounding at each reasoning step through two sequential phases: (1) Verification, which performs object extraction and quantifies chain-level faithfulness ($F_{\text{chain}}$), and (2) Action based on a pre-defined faithfulness threshold. The framework is realized via a general, lightweight, and extensible interface composed of callable functions: Poll(), Ground(), Select(), Abstain(), and Count().
  • Extensive experiments across multiple benchmarks reveal that current MLLMs (Qwen-2.5-VL-7B, InternVL3-8B, and LLaVA-One-Vision-1.5-8B) consistently overestimate their perceptual grounding. Our results show that enforcing stepwise verification is crucial for enhancing perceptual without compromising task accuracy. We comprehensively benchmark against two common prompt-based (CoT, VAT) and tool-augmented (Grounded-CoT, ReAct) baselines, demonstrating that FaithAct achieves the highest overall perceptual faithfulness, providing empirical evidence of its ability to mitigate hallucination.
2 Related Work
2.1 Reasoning Frameworks for MLLMs
MLLMs have developed surprising reasoning capabilities in multimodal problem solving, task planning, scientific discovery and so on lumathvista; li2025chemvlm; gao2024cantor; an2024mc; an2025unictokens; lin2025perceive; li2025miv; li2025taco. To enhance systematic reasoning, several frameworks have been proposed to decompose multimodal problems into interpretable steps. Chain-of-Thought (CoT) chen2024m3cot; sun2025mm; shao2024visual; zhou2025logic; wang2025vl encourages models to generate intermediate steps before producing final answers, improving step-by-step correctness. Grounded-CoT Wu2025GroundedCF recently extends this idea by explicitly linking each reasoning step to the grounding coordinates, but faithfulness was not conidered, for example models behave unfaithfully, while our work aims to enforce evidence before reasoning continues. ReAct-style frameworks yao2022react further integrate reasoning with environment interaction, allowing models to iteratively query or retrieve external information. More recently, Visual Abstract Thinking (VAT) liu2025visual explores hierarchical and compositional reasoning, encouraging models to form abstract visual concepts that support complex decision-making. Such generated reasoning chains, though often fluent and logically structured, may still include steps unsupported by visual evidence or inconsistent with the model’s actual decision process yu2024rlhf. More importantly, prior work has rarely incorporated faithfulness as an explicit design principle, leaving open the challenge of ensuring that each intermediate step remains perceptually grounded in the input before the next step continues.
2.2 Faithfulness in Multimodal Reasoning
Despite the demonstrated effectiveness of reasoning in improving the performance of MLLMs, an unfaithful problem emerges: the reasoning traces generated by these models do not always behave faithfully to the internal processes that produce their final answers lyu2023faithful; turpin2023language; lanham2023measuring. Models often rely on latent knowledge or shortcut associations that are not explicitly expressed in their reasoning chains. Empirical analyses have shown that such traces can be post-hoc rationalizations, fluent but unfaithful narratives, shortcut reasoning, or approximate reconstructions of what the model actually “thinks”, rather than faithful records of how the model arrived at its outputs. As a result, the generated reasoning steps may read as a plausible but untrustworthy explanation, motivating the need for explicit faithfulness assessment jacovi2020towards; parcalabescu2023measuring; matton2025walk; barez2025chain; arcuschin2025chain. Several works have introduced benchmarks or metrics to evaluate faithfulness explicitly. M4 li2023m4, FaithScore jing2023faithscore and TIFA hu2023tifa propose metrics for evaluating the faithfulness for vision-language models. However, our findings reveal that behavioral alignment (behavioral faithfully) does not guarantee the correctness of the final output (see left panel in Fig. 1).
Object Hallucination as Unfaithful Consequences.
Object hallucination is identified as a common challenge, as a consequence, in large vision-language models (LVLMs), where the model describes or reasons about objects absent in the input image. Several benchmarks, such as POPE Li2023EvaluatingOH and MMHal-Bench sun2024aligning, have been developed to systematically evaluate this phenomenon. However, these efforts typically treat hallucination as an isolated failure mode rather than as a manifestation of broader unfaithful reasoning. In contrast, we view hallucination as one instance of perceptual unfaithfulness, and hallucination can be mitigated by explicitly grounding objects within intermediate reasoning steps.
Recent studies address hallucination through different mechanisms: (i) training-time alignment, such as hallucination-aware RLHF or preference optimization sun2024aligning; zhang2024hallucination; (ii) decoding-time constraints, including grounded or contrastive decoding Li2023EvaluatingOH; and (iii) feature-level grounding strategies that enhance cross-modal alignment ghosh2024vdgd. Together, these advances underscore that faithful reasoning requires not only linguistic coherence but also perceptual accountability, further motivating to verify evidential grounding before inference.
@@ 3 Perceptual and Behavioral Faithfulness
Building on our observation above, we have classified two “unfaithful” cases, and we further discuss these two and our focus in this section.
  • Perceptual faithfulness (PF) concerns whether each reasoning step is grounded in verifiable input evidence, i.e., whether the model “sees” only what is perceptually present in the image or context.
  • Behavioral faithfulness (BF) concerns whether the reasoning trace accurately reflects the internal decision pathway that leads to the model’s final output [^1] Noted that the model is not required to output the correct answer (jacovi2020towards; dasgupta2022framework)..
Intuitively, perceptual faithfulness asks “Does the model reason only from what it actually perceives?” while behavioral faithfulness asks “Does the reasoning trace truly reflect how the model arrived at its answer?” Although both are desirable, they serve different purposes: perceptual faithfulness ensures grounding and verifiability, while behavioral faithfulness ensures introspective consistency.
@@
Our focus on perceptual faithfulness is motivated by the hypothesis that when reasoning remains perceptually grounded, behavioral consistency tends to follow. Practically, perceptual faithfulness offers a more actionable and operationally measurable objective, each reasoning step can be directly validated against the available multimodal evidence by inference. In contrast, behavioral faithfulness depends on inaccessible internal dynamics of large models and is therefore treated as a consequence rather than a controllable design target. We leave direct evaluation of behavioral faithfulness for future work.
4 FaithEval: Perceptual Faithfulness Evaluation Pipeline
Figure 2: Overview of the proposed FaithAct framework. FaithAct enforces faithfulness-first multimodal reasoning through two sequential phases, namely verification and action. Left: The Quantifying Faithfulness stage (Sec. 4) extracts claimed objects from each reasoning step and verifies their perceptual grounding in the image using callable functions such as Poll() and Ground(). Right: The Faithful-First Planning stage (Sec. 5) refines reasoning by selectively retaining only evidence-supported objects through structured APIs (Select(), Abstain(), Count()). This design ensures faithfulness as a principle, where every reasoning step is evidentially justified before contributing to the final answer.
@@ 4 FaithEval: Perceptual Faithfulness Evaluation Pipeline
We operationalize perceptual faithfulness by first evaluating how well each reasoning step aligns with verifiable evidence in the input. To this end, we design a general-purpose pipeline that systematically evaluates the degree of perceptual grounding across reasoning steps. The pipeline consists of three key stages:
  1. (i)
    Extraction of Claimed Objects: identifying all visual entities or attributes referenced in each reasoning step;
  2. (ii)
    Preference Polling and Grounding: retrieving candidate evidence from the input (e.g., detected regions, captions, or image–text alignments) and associating them with the corresponding claims;
  3. (iii)
    Faithfulness Scoring: quantifying how faithfully each reasoning step and the entire reasoning chain aligns with perceptual evidence in the input.
@@ 4.1 Extraction of Claimed Objects
We begin by defining the setting for perceptual faithfulness verification. Given a multimodal input consisting of an image $I$ and and a corresponding textual query $Q$, an MLLM produces a reasoning chain by prompting “think step by step” wei2022chain.
$R_{raw}=\{s_{1},s_{2},\dots,s_{T}\},$ (1)
where each $s_{t}$ represents an intermediate reasoning step generated by the model. To quantify perceptual faithfulness, we must identify all claimed objects across both the question and the reasoning steps.
We therefore employ a structured extraction process to isolate visually grounded claims from textual noise. Specifically, we process the concatenated text of $Q$ and each reasoning step $s_{t}$ using a helper LLM to extract meaningful object mentions. Each reasoning step is independently analyzed, excluding introductory or concluding phrases (e.g., “Let’s begin reasoning” or “Therefore, the answer is…”) to focus solely on evidence-bearing content. Concretely, we query Qwen2.5-7B-Instruct qwen2.5 with a structured prompt (see Appendix A) to extract, for each step $s_{t}$, a set $O_{t}$ of $m$ claimed objects:
$O_{t}=\{O_{t}^{1},O_{t}^{2},\dots,O_{t}^{m_{t}}\}.$ (2)
The union of all extracted objects across the reasoning chain is denoted as
@@ 4.2.1 Preference Polling
While object detectors can localize objects spatially, their confidence scores are often unreliable when visual cues are weak, such as in cases of occlusion, clutter, or low salience. To address this limitation, we introduce a lightweight preference polling module that performs an initial verification of object existence prior to grounding. This module serves as an evidence gate, polling the visual scene to estimate whether an object is perceptually verifiable at all. By filtering out unsupported object claims early, the polling stage ensures that only objects with sufficient perceptual evidence proceed to subsequent grounding and faithfulness scoring.
Method.
We employ a frozen CLIP-ViT-Large model radford2021learning as the multimodal encoder pair $(f_{\text{img}},f_{\text{txt}})$ and attach a lightweight polling head, a two-layer MLP with GELU activation, to predict whether a given object exists in the image. The polling model is trained on the POPE dataset Li2023EvaluatingOH, which provides binary image–object existence labels (see Appendix B for details).
@@
where $c_{p,t}^{i}$ corresponds to the predicted existence probability for each claimed object $O_{t}^{i}$. Objects whose confidence falls below a threshold $\tau_{p}$ are considered unverified and excluded from subsequent grounding. This early verification step thus acts as a perceptual filter, aligning with our evidence-first philosophy.
@@ 4.2.2 Grounding
After preference polling confirms the likely existence of an object, we proceed to localize it within the image using a grounded detection model. This step provides region-level visual evidence and confidence scores for the subset of objects verified by polling.
Method.
For each verified object $O_{t}^{i}$ (where $c_{p,t}^{i}>\tau_{p}$), we employ a frozen GroundingDINO-base model liu2023grounding to detect and ground the object in the input image $I$. GroundingDINO returns a set of bounding boxes and associated confidence scores:
@@
While preference polling provides global existence verification, grounding supplies fine-grained spatial evidence. Combining these complementary signals yields a more reliable perceptual faithfulness measurements.
@@ 4.3 Faithfulness Scoring
5 FaithAct: Faithfulness-First Planning
Having established how to quantify perceptual faithfulness, we now turn to the problem of how to integrate faithfulness into the reasoning process itself. We propose FaithAct, a Faithfulness-First Planning framework that enforces evidential verification during reasoning generation. Unlike conventional generate-then-verify paradigm, FaithAct follows a verify-as-you-generate principle, where each step is explicitly checked for perceptual support before being admitted into the reasoning chain.

### v3

# Faithful-First Reasoning, Planning, and Acting for Multimodal LLMs

@@ Abstract
Multimodal Large Language Models (MLLMs) frequently suffer from unfaithfulness, generating reasoning chains that drift from visual evidence or contradict final predictions. We propose Faithful-First Reasoning, Planning, and Acting (RPA) framework in which FaithEvi provides step-wise and chain-level supervision by evaluating the faithfulness of intermediate reasoning, and FaithAct uses these signals to plan and execute faithfulness-aware actions during inference. Experiments across multiple multimodal reasoning benchmarks show that faithful-first RPA improves perceptual faithfulness by up to 24% over prompt-based and tool-augmented reasoning frameworks, without degrading task accuracy. Our analysis shows that treating faithfulness as a guiding principle perceptually faithful reasoning trajectories and mitigates hallucination behavior. This work thereby establishes a unified framework for both evaluating and enforcing faithfulness in multimodal reasoning. Code will be released upon acceptance.
Faithful-First Reasoning, Planning, and Acting for Multimodal LLMs
Junxian Li1thanks: Equal contribution.   Xinyue Xu211footnotemark: 1   Sai Ma311footnotemark: 1   Di Zhang4   Sichao Li5thanks: Correspondence: scli@cityu.edu.mo 1Shanghai Jiao Tong University 2The Hong Kong University of Science and Technology 3The Australian National University 4 Fudan University 5City University of Macau
@@ 1 Introduction
Despite rapid progress in multimodal large language models (MLLMs) Sun et al. (2023); Chen et al. (2024b); An et al. (2025b); Bai et al. (2025), their reasoning traces remain unfaithful: models frequently produce persuasive explanations that conflict with perceptual evidence, or utilize post-hoc rationalizations fabricate their reasoning progress Arcuschin et al. (2025); Barez et al. (2025). This gap poses a central challenge for trustworthy reasoning. Existing efforts typically focus on improving task accuracy or enriching Chain-of-Thought (CoT) generation Zhang et al. (2023), yet the unfaithful reasoning remains unaddressed.
@@
We are motivated by the following principle and observations Goyal et al. (2017).
@@
This principle echoes long-standing findings in Visual Question Answering (VQA): systems should avoid answering beyond available evidence and resist over-reliance on language priors Antol et al. (2015); Agrawal et al. (2018); Bender et al. (2021).
Figure 1: Perceptually and behaviorally unfaithful examples. Left: Perceptually unfaithful and behaviorally unfaithful. The model presents a step-by-step reasoning trace describing alternating arrow directions and increasing dot counts, yet such reasoning does not reflect its actual decision process. The final choice (A) is likely made through pattern association, with the explanation generated post hoc to rationalize it. Right: Perceptually unfaithful but behaviorally faithful. The reasoning aligns with its final answer but incorrectly describes the bicycle as yellow, influenced by a nearby yellow bus. This illustrates a visually plausible but perception-unfaithful reasoning step, where linguistic association overrides perceptual grounding.
As illustrated by examples in Fig. 1, we observe that MLLMs can generate plausible explanations that are perceptually inconsistent with the underlying visual input, regardless of whether the final prediction is correct. To formalize this observation, we distinguish perceptual faithfulness (reasoning steps align with the model’s input) from the predominant focus in prior work on behavioral faithfulness (reasoning steps align with the model’s output) Arcuschin et al. (2025); Matton et al. (2025); Ming et al. (2024); Li et al. (2023b).
Our answer. We argue that faithfulness should be a design principle, not merely a post-hoc evaluation objective. Reasoning frameworks should explicitly verify the evidential grounding of each step before it is admitted into the chain, ensuring that reasoning remains both perceptually grounded and behaviorally aligned.
@@
This motivates the Faithful-First Reasoning, Planning, and Acting (RPA) framework, which operationalizes the principle as a design constraint, enforcing faithfulness throughout multimodal reasoning. Within this framework, we make the following contributions:
@@
$\bullet$ We introduce FaithEvi, an evidence-based perceptual faithfulness evaluation pipeline via evidence that extracts claimed objects from each reasoning step and verifies their existence at both local and global levels through preference polling and visual grounding. It assigns a step-level faithfulness score ($F_{\text{step},t}$) to each reasoning step and aggregates them into a chain-level score ($F_{\text{chain}}$), providing a general quantitative measure of perceptual faithfulness with respect to the input evidence.
$\bullet$ We propose FaithAct, a faithful-first planning and acting mechanism that enforces evidential grounding during inference. It operates through two sequential phases: (1) Planning, which leverages FaithEvi signals to assess the current chain’s faithfulness ($F_{\emph{chain}}$) and determine admissible next steps; and (2) Acting, which executes faithfulness-aware actions subject to dynamic thresholds. The mechanism is realized via a lightweight, extensible interface of callable functions.
$\bullet$ Extensive experiments across multiple benchmarks reveal that current MLLMs consistently underestimate evidential support in reasoning process. We demonstrate that enforcing step-wise verification significantly enhances perceptual faithfulness without compromising task accuracy. Compared to standard prompt-based and tool-augmented frameworks, Faithful-First RPA achieves the highest overall perceptual faithfulness and effectively mitigates content hallucination.
@@ 2 Perceptual and Behavioral Faithfulness
Building on the motivating observations (Fig. 1) and prior analyses of explanation faithfulness (Appx. A), we distinguish two “unfaithful” notions and clarify our focus of this work.
Previous studies mainly concern whether a reasoning trace accurately reflects the decision process that led to the model’s final output Lyu et al. (2023); Matton et al. (2025); Barez et al. (2025); Arcuschin et al. (2025), which we refer to as Behavioral faithfulness (BF). While important, it underestimates the importance of input evidence in the reasoning process. We therefore propose Perceptual faithfulness (PF), which concerns whether individual reasoning steps are grounded in verifiable input evidence, such as entities and attributes that are perceptually present in the input. A perceptually unfaithful explanation may invoke objects and properties that are unsupported or contradicted by the visual input. Importantly, the model is not required to output the correct answer to be faithful (Jacovi and Goldberg, 2020; Dasgupta et al., 2022).
For instance, the left example in Fig. 1 shows a correct prediction accompanied by reasoning that is neither perceptually grounded nor behaviorally aligned. The right example illustrates reasoning that aligns with the model’s prediction but remains perceptually unfaithful, as it relies on visually unsupported attributes (e.g., bicycle color). While ideal reasoning would be both perceptually and behaviorally faithful and yield a correct prediction, optimizing all three remains challenging. In this work, we focus on perceptual faithfulness.
@@
This choice is motivated by the hypothesis that when reasoning remains perceptually grounded, behavioral consistency tends to follow, without introducing additional optimization objectives such as output correctness. Practically, perceptual faithfulness offers a more actionable and operationally measurable objective, each reasoning step can be directly validated against the available multimodal evidence by inference. In contrast, behavioral faithfulness depends on inaccessible internal dynamics of MLLMs and is therefore treated as a consequence rather than a controllable design target.
3 FaithEvi: Perceptual Faithfulness Evaluation Pipeline
Figure 2: Faithful-first reasoning, planning, and acting framework. Given an image-question pair, FaithEvi evaluates the perceptual faithfulness of intermediate reasoning, producing step- and chain-level faithfulness scores. Guided by these signals, FaithAct plans and acts faithfulness-aware actions during inference.
@@ 3 FaithEvi: Perceptual Faithfulness Evaluation Pipeline
We operationalize perceptual faithfulness by first evaluating how well each reasoning step aligns with the input verifiable evidence. To this end, we design a general-purpose pipeline that systematically evaluates the degree of perceptual grounding across reasoning process at step-wise and chain-level. The pipeline consists of three key stages: (i) Extraction of Claimed Objects (ii) Preference Polling and Grounding (iii) Faithfulness Scoring.
@@ 3.1 Extraction of Claimed Objects
We begin by defining the setting for perceptual faithfulness verification. Given a multimodal input consisting of an image $I$ and and a corresponding textual query $Q$, an MLLM produces a reasoning chain by prompting “think step by step” Wei et al. (2022): $R_{raw}=\{s_{1},s_{2},\dots,s_{T}\},$ where each $s_{t}$ represents an intermediate reasoning step generated by the model. To quantify perceptual faithfulness, we must identify all claimed objects across both the question and the reasoning steps.
We therefore employ a structured extraction process to isolate visually grounded claims from textual noise. Specifically, we process the concatenated text of $Q$ and each reasoning step $s_{t}$ using a helper LLM to extract meaningful object mentions. Each reasoning step is independently analyzed, excluding introductory or concluding phrases (e.g., “Let’s begin reasoning” or “Therefore, the answer is…”) to focus solely on evidence-bearing content.
Concretely, we query Qwen2.5-7B-Instruct (Qwen)  Team (2024) with a structured prompt (see Appx. B) to extract, for each step $s_{t}$, a set $O_{t}$ of $m_{t}$ claimed objects, where $O_{t}=\{O_{t}^{1},O_{t}^{2},\dots,O_{t}^{m_{t}}\}.$ The union of all extracted objects across the reasoning chain is denoted as
@@ 3.2.1 Preference Polling
While object detectors can localize objects spatially, their confidence scores are often unreliable when visual cues are weak, such as in cases of occlusion, clutter, or low salience. To address this limitation, we introduce a lightweight preference polling module that performs an initial verification of object existence prior to grounding. This module serves as an evidence gate, polling the visual scene to estimate whether an object is perceptually verifiable.
We employ a frozen CLIP-ViT-Large (CLIP) model Radford et al. (2021) as the multimodal encoder pair $(f_{\text{img}},f_{\text{txt}})$ and attach a lightweight polling head, a two-layer MLP with GELU activation, to predict whether a given object exists in the image. The polling model is trained on the POPE dataset Li et al. (2023b), which provides binary image–object existence labels (see Appx. C).
@@
where $c_{p,t}^{i}$ corresponds to the predicted existence probability for each claimed object $O_{t}^{i}$. This early verification step thus acts as a perceptual verifier, aligning with our evidence-first philosophy.
@@ 3.2.2 Grounding
After preference polling confirms the likely existence of an object, we proceed to localize it within the image using a grounded detection model. This step provides region-level visual evidence and confidence scores for objects.
For each object $O_{t}^{i}$ we employ a frozen GroundingDINO-base model Liu et al. (2023b) to detect and ground the object in the input image $I$. GroundingDINO returns a set of bounding boxes and associated confidence scores:
@@
While preference polling provides global existence verification, grounding supplies fine-grained spatial evidence. Combining these complementary signals yields a more reliable perceptual faithfulness measurement.
@@ 3.3 Faithfulness Scoring
4 FaithAct: Faithful-First Planning and Acting
Having established how to quantify perceptual faithfulness, we now turn to the problem of how to integrate faithfulness into the reasoning process itself. We propose FaithAct, a Faithfulness-First Planning and Acting framework that enforces evidential verification during reasoning generation. Unlike conventional generate-then-verify paradigm, FaithAct follows a verify-as-you-generate principle, where each step is explicitly checked for perceptual support before being admitted into the reasoning chain.